The ratio of the complement of an angle to the supplement of an angle is 18.5 to 41. Find the measure of the complement.
Answer by amarjeeth123(570)

Let the angle be x degrees.
Complement of the angle=90-x
Supplement of the angle=180-x
(90-x)/(180-x)=18.5/41
41(90-x)=18.5(180-x)
(41-18.5)x=3690-3330
22.5x=360
x=16
The complement of the angle is (90-16) i.e.74 degrees.